# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision Language Models

### v2

@@ 4.2.1 Recognition
Table 2 shows the Recognition Score of models on Description task. GPT-4V achieves the best performance in terms of recognition, which means it can recognize and describe more entities. Besides, it is significantly better than other open-source LVLMs, indicating open-source LVLMs still have some room for development before reaching the recognition capability of GPT-4V. ShareGPT4V-7B and ShareGPT4V-13B achieves better performance than other open-source LVLMs. As it follows the design of LLaVA v1.5, one possible reason is that ShareGPT4V uses a high-quality image-text dataset featuring highly descriptive captions for training, which makes it describe more entities shown in images. Though GPT-4V achieves the best performance, there are still a lot of entities that are not recognized by GPT-4V, indicating a room for improvement.
@@
InstructBLIP-7B 0.50
Qwen-VL-Chat 0.54
LLaVA-v1.5-7B 0.51
LLaVA-v1.5-13B 0.52
mPLUG-Owl-2 0.48
ShareGPT4V-7B 0.57
ShareGPT4V-13B 0.60
GPT-4V 0.77
Human 0.94
Table 2: Recognition score of LVLMs on Description task. For reference, the Recognition Score of Human is calculated based on the annotated [Description] in CogBench as an estimate of an upper limit score.
@@ 4.2.2 Cognition
InstructBLIP-7B 0.13 0.54 0.19 0.35 0.10 0.04 0.05 0.17 0.17
Qwen-VL-Chat 0.26 0.54 0.31 0.28 0.18 0.13 0.09 0.20 0.22
LLaVA-v1.5-7B 0.09 0.52 0.19 0.27 0.09 0.05 0.05 0.19 0.16
LLaVA-v1.5-13B 0.04 0.47 0.21 0.33 0.09 0.06 0.04 0.20 0.17
mPLUG-Owl-2 0.04 0.47 0.19 0.28 0.10 0.05 0.05 0.14 0.15
ShareGPT4V-7B 0.22 0.56 0.21 0.25 0.09 0.03 0.05 0.19 0.16
ShareGPT4V-13B 0.22 0.54 0.24 0.29 0.09 0.06 0.09 0.17 0.17
GPT-4V 0.44 0.71 0.45 0.29 0.36 0.37 0.18 0.54 0.41
Human 0.87 0.95 0.95 0.84 0.99 0.94 0.86 0.91 0.93
Table 3: Cognition Scores of LVLMs on Description task evaluated by GPT-4. For reference, the Cognition Score of Human is calculated based on the annotated [Description] in CogBench as an estimate of an upper limit score.
Table 3 shows Cognition Scores of LVLMs on Description task. Similarly, GPT-4V achieves the best performance and there is a large performance gap between GPT-4V and other open-source models. For open-source models, Qwen-VL-Chat achieves the best performance. Though ShareGPT4V achieves better recognition performance than other open-source LVLMs, it’s cognition performance does not show a significant improvement compared with other models. In terms of different reasoning capabilities, all of the LVLMs show better performance on [Location Reasoning] than others, which is probably because it is a kind of relatively lower-level reasoning. Differently, for [Event Reasoning], [Event Relationship Reasoning], and [Next Moment Event Reasoning], all of the open-source LVLMs show very low performance, indicating they almost do not understand the story in the images at all. The significantly large performance gap on these three kinds of reasoning types between GPT-4V and open-source LVLMs could be a manifestation of the capabilities emerging in GPT-4V. Though GPT-4V achieves the best performance, there is also a huge gap between the cognitive ability of LVLMs and human beings, and the gap is obviously larger than in the aspect of recognition. This indicates that LVLMs still have a lot of room for development in terms of cognitive abilities.
@@ 4.2.3 Case Study of Description Task
Figure 5 shows a case of Description task. In terms of recognition, GPT-4V shows a good performance by recognizing most annotated entities (“boy”, “girl”, “man”, “woman”, “table”, “dress”, “hat”, “suit”) and only fails to recognize the “sofa”, “television”, “handbag”. LLaVA-v1.5-13B obviously recognizes fewer entities than GPT-4V, it only recognizes “boy”, “woman”, and “dress”. However, GPT-4V fails to understand the story in the image and gets a 0 in terms of cognition. This is because it neither clarifies the location nor the relationships between the characters, nor does it accurately depict the events occurring in the image, etc. LLaVA-v1.5-13B is slightly better and successfully inferred that the woman is the mother.
The case shows CogBench can reflect that current LVLMs fail at aspects like recognition and cognition and the cognitive abilities exhibited in the description still have some gap with the level of humans.
@@ 4.3 Results of VQA Task
InstructBLIP-7B 0.34 0.57 0.47 0.52 0.36 0.44 0.39 0.37 0.43
Qwen-VL-Chat 0.64 0.84 0.63 0.55 0.54 0.40 0.50 0.53 0.55
LLaVA-V1.5-7B 0.55 0.80 0.54 0.54 0.51 0.37 0.47 0.53 0.52
LLaVA-V1.5-13B 0.68 0.85 0.60 0.62 0.48 0.54 0.68 0.56 0.59
mPLUG-Owl-2 0.57 0.82 0.54 0.53 0.51 0.46 0.53 0.56 0.55
ShareGPT4V-7B 0.62 0.78 0.63 0.59 0.50 0.36 0.53 0.56 0.54
ShareGPT4V-13B 0.70 0.85 0.59 0.56 0.50 0.50 0.72 0.58 0.58
GPT-4V 0.72 0.88 0.79 0.69 0.68 0.65 0.72 0.71 0.71
Human 0.98 0.96 0.99 0.96 0.97 0.98 0.96 0.95 0.96
Table 4 shows the performance of LVLMs on VQA task. Consistent with results in previous sections, GPT-4V achieves the best performance and there is a performance gap between GPT-4V and open-source models. For open-source models, ShareGPT4V-13B and LLaVA-v1.5-13B achieves better performance than other LVLMs based on 7B LLMs, which shows the importance of LLM size to LVLMs. As for LVLMs based on 7B LLMs, the performance of InstructBLIP is the worst and there is also a gap between InstructBLIP and other models. One possible reason is that it only finetunes the Q-former for instruction-tuning as Q-former has limited capacity compared with LLMs. The performance of other models are similar. Consistent with previous findings, reasoning about location is also the easiest for LVLMs and event-related reasoning are more difficult. There is also a large gap between the performance of LVLMs and human beings. Note that the accuracy of Human in Table 4 is calculated based on the responses of a healthy 24-year-old male with a Bachelor’s degree.
@@ 5 Related Work
To better understand emerging capabilities of LVLMs, many scholars are dedicated to proposing different evaluation benchmarks. LVLM-eHub Xu et al. (2023) evaluates LVLMs’ multimodal capabilities across six categories using various publicly available computer vision datasets. MME Fu et al. (2023), MMBench Liu et al. (2023c) and SEED Bench Li et al. (2023a) use True/False Questions or Multiple-Choice Questions to evaluate different abilities of LVLMs. MM-VET Yu et al. (2023) evaluates LVLMs in terms of their integrated VL capabilities via open-ended questions. Different from them, CogBench mainly focus on high-level cognition evaluation of LVLMs. Though some of them also consider cognition as one of the evaluation dimensions, most images they use can only be used to evaluate limited aspects of reasoning.
@@
Image Captioning is a classical VL task. As model capabilities advance, researchers strive to enhance their ability to describe images in detail. Krause et al. (2017) propose Image Paragraph Captioning, tasking models with generating a descriptive paragraph for an image. Recently, some researchers Xie et al. (2022); Zhu et al. (2023a); Zhuge et al. (2023); Chen et al. (2023a) are trying to leveraging the ability of LLMs to generate more detailed image descriptions. None of these tasks consider to evaluate the high-level cognitive ability of models through description. HL dataset Cafagna et al. (2023) requires models to generate high-level captions, but it only considers three aspects (scene, action, rationale). The content of most images in existing datasets does not reach the level of a story. This reveals the need for higher-level datasets like CogBench.
@@
Visual Reasoning is closely related to the cognitive ability of models. VCR Zellers et al. (2019) tasks models with answering visual questions using commonsense reasoning and justifying their answers. VisualCOMET Park et al. (2020) is a framework of visual commonsense reasoning tasks to predict past, future events, and present intents. Hessel et al. (2022) utilize images from VCR and Visual Genome Krishna et al. (2017) to evaluate the ability of models to perform abductive reasoning. Fu et al. (2022) propose a task to identify the time and location of a given image. CURE Chen et al. (2023b) is proposed to measure both the zero-shot reasoning performance and consistency of VLMs. Similarly, compared CogBench, these tasks considers less kinds of reasoning and CogBench can be seen as the next step of these efforts.
@@ 6 Conclusion
Limitations
Compared to some large-scale benchmarks, CogBench is a relatively light-weight benchmark with fewer images due to our strict image collection criteria. We aim to substitute quantity with high-quality images and annotations containing rich semantic information. In the future, we will keep updating CogBench to include more high-quality images and annotations.
Ethical Considerations
Most images in CogBench are manually collected from Pinterest. We follow the terms of service of Pinterest to collect these images. The images are used as fair use for research purposes only. We will share the data with other researchers who will follow the ethical considerations as established in this study. As for annotation, we pay our annotators a fair wage and will answer any questions they have during their annotation process promptly. They are free to take breaks or quit the annotation task at any time.
@@ Appendix C Prompt of GPT-based Description Cognition Evaluation
Given a <DESCRIPTION> and some <KEY POINT>s, please tell me if the <DESCRIPTION> explicitly presents the exact or similar semantics of each <KEY POINT>. The following points are required: 1) Instead of reasoning about whether each <KEY POINT> is possibly correct based on the <DESCRIPTION>, you only need to determine whether the <DESCRIPTION> mentions the semantics in the <KEY POINT>. 2) Do not overlook the semantics in the <DESCRIPTION> that are semantically equivalent to the <KEY POINT> but expressed in different ways. For instance, if the <DESCRIPTION> mentions "The woman is playing with her son…", we can tell it successfully includes semantics in the <KEY POINT> "The woman is the mother of the boy." 3) If several possible scenarios are listed using ’or’ at a <KEY POINT>, you only need to determine whether one of these scenarios is mentioned in the <DESCRIPTION>. Assign a score of 0 or 1 to each <KEY POINT>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by models.} <KEY POINT>: 1. {Annotated key point 1.} 2. {Annotated key point 2.} N. {Annotated key point N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <KEY POINT>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 8: Evaluation prompt for GPT-4 of Reseasoning types other than [Event Relationship Reasoning].
@@
Figure 9: Evaluation prompt for GPT-4 of [Event Relationship Reseasoning].
@@ Appendix D Introduction to Selected LVLMs
  • InstructBLIP Dai et al. (2023) builds upon BLIP-2 Li et al. (2023b). It consists of an image encoder, a LLM, and a Query Transformer (Q-Former). During instruction tuning, only the Q-Former is updated. We use “blip2-vicuna-instruct” for testing.
  • Qwen-VL-Chat Bai et al. (2023) is a instruction-tuned VL chatbot based on Qwen-VL. Its training process consists of two stages of pre-training and a final stage of instruction fine-tuning. As for architecture, it consists of a vision encoder, a LLM, and position-aware vision-language adapter. We test “Qwen-VL-Chat”.
  • LLaVA v1.5 (Liu et al., 2023a) is an upgraded version of LLaVA Liu et al. (2023b). LLaVA connects a vision encoder and LLM for general-purpose visual and language understanding. It is instruction-tuned on the language-image instruction-following data generated by language-only GPT-4 OpenAI (2023). By using CLIP-ViT-L-336px with an MLP projection and adding academic-task-oriented VQA data with simple response formatting prompts, LLaVA v1.5 achieves better performance. “llava-v1.5-7b” and “llava-v1.5-13b” are tested.
  • mPLUG-Owl-2 Ye et al. (2023) mainly comprises a fundamental vision encoder, a visual abstractor, and a language decoder. It also adopts a two-stage training strategy, comprising pre-training and visual instruction tuning. We test “mplug-owl2-llama2-7b”.
  • ShareGPT4V Chen et al. (2023a) follows the design of LLaVA v1.5. They incorporate a large-scale resource featuring highly descriptive captions into both the pre-training and supervised fine-tuning phases of ShareGPT4V model. We test“ShareGPT4V-7B” and “ShareGPT4V-13B”.
  • GPT-4V (OpenAI, 2023) is one of the most powerful LVLMs in the world developed by OpenAI. The version of “gpt-4-vision-preview” is tested.
Table 5 shows an overview of the design of different LVLMs.
@@
Table 5: LVLMs evaluated in this paper.

### v3

@@ 3.3.1 Recognition
Table 2: Recognition score of LVLMs on Description task. For reference, the Recognition Score of Oracle is calculated based on the annotated [Description] in CogBench as an estimated upper bound.
@@
InstructBLIP-7B 0.40
Qwen-VL-Chat 0.42
LLaVA-v1.5-7B 0.40
LLaVA-v1.5-13B 0.41
mPLUG-Owl-2 0.37
ShareGPT4V-7B 0.47
ShareGPT4V-13B 0.49
GPT-4V 0.67
GPT-4o 0.72
Oracle 0.93
Table 2 shows the Recognition Score of models on Description task. GPT-4o achieves the best performance in terms of recognition, which means it can recognize and describe more entities. Besides, GPT-4o and GPT-4V are significantly better than other open-source LVLMs, indicating these open-source LVLMs still have some room for development before reaching the recognition capability of GPT-4. ShareGPT4V-7B and ShareGPT4V-13B achieves better performance than other open-source LVLMs. As it follows the design of LLaVA v1.5, one possible reason is that ShareGPT4V uses a high-quality image-text dataset featuring highly descriptive captions for training, which makes it describe more entities shown in images. GPT-4o, while top-performing, still misses many entities, suggesting room for improvement even on recognition capability.
@@ 3.3.2 Cognition
Table 3 shows Cognition Scores of LVLMs on Description task. Similarly, GPT-4o achieves the best performance and there is a large performance gap between GPT-4 and other open-source models. For open-source models, Qwen-VL-Chat achieves the best performance. In terms of different reasoning capabilities, all of the LVLMs show better performance on [Location Reasoning] than others, which is probably because it is a kind of relatively lower-level reasoning. Differently, for [Event Reasoning], [Event Relationship Reasoning], and [Next Moment Event Reasoning], all of the open-source LVLMs show very low performance, indicating they almost do not understand the story in the images at all. The significantly large performance gap on these three kinds of reasoning types between GPT-4 and open-source LVLMs could be a manifestation of the capabilities emerging in GPT-4. Though GPT-4o achieves the best performance, there is also a huge gap between its ability and the Oracle, and the gap is obviously larger than that of recognition scores. This indicates that LVLMs still have a lot of room for development in terms of cognitive abilities.
Table 3: Cognition Scores of LVLMs on Description task evaluated by GPT-4. For reference, the Cognition Score of Oracle is calculated based on the annotated [Description] in CogBench as an estimated upper bound.
@@
InstructBLIP-7B 0.15 0.54 0.25 0.30 0.10 0.05 0.02 0.17 0.17
Qwen-VL-Chat 0.28 0.57 0.28 0.26 0.14 0.12 0.04 0.13 0.19
LLaVA-v1.5-7B 0.09 0.45 0.15 0.18 0.09 0.05 0.02 0.13 0.13
LLaVA-v1.5-13B 0.13 0.49 0.18 0.24 0.10 0.05 0.04 0.16 0.15
mPLUG-Owl-2 0.06 0.48 0.24 0.21 0.08 0.04 0.02 0.14 0.14
ShareGPT4V-7B 0.19 0.60 0.21 0.22 0.10 0.04 0.03 0.15 0.16
ShareGPT4V-13B 0.23 0.57 0.24 0.26 0.12 0.07 0.03 0.14 0.17
GPT-4V 0.43 0.71 0.47 0.24 0.35 0.31 0.10 0.46 0.37
GPT-4o 0.51 0.77 0.55 0.37 0.44 0.39 0.15 0.40 0.43
Oracle 0.92 0.98 0.94 0.81 0.98 0.92 0.90 0.92 0.93
@@ 3.3.3 Case Study of Description Task
Figure 5 shows a case of Description task. In terms of recognition, GPT-4V shows a good performance by recognizing most annotated entities such as boy, girl, man, woman, table, dress, hat, suit and only fails to recognize sofa, television, handbag . LLaVA-v1.5-13B obviously recognizes fewer entities than GPT-4V, as it only recognizes boy, woman, dress. However, GPT-4V fails to understand the story in the image and gets a 0 in terms of cognition. This is because it neither clarifies the location nor the relationships between the characters, nor does it accurately depict the events occurring in the image, etc. LLaVA-v1.5-13B is slightly better and successfully inferred that the woman is the mother. This case demonstrates that CogBench reveals current LVLMs falling short in recognition and cognition, with a gap remaining between their cognitive abilities and human levels.
@@
3.3.4 Effectiveness Analysis of GPT-based Cognition Evaluation of Description Task
Table 4: CoR accuracy of evaluation methods.
Model Accuracy
ROUGE 0.656
BERTScore 0.635
BLEURT 0.620
DeBERTa 0.693
DocNLI 0.714
GPT-3.5 0.807
GPT-4 0.833
To validate GPT-based cognition evaluation, we manually scored the CoRs of 20 images with a binary scale (0/1) and compared various evaluation methods’ accuracy on this subset. Table 4 reveals that GPT-4 offers the highest accuracy, demonstrating that GPT-based evaluation aligns well with human assessment. Therefore it is effective to assess LVLMs’ performance on the Description task. Implementation details of evaluation methods beyond ChatGPT and GPT-4 can be found in Appendix F.
@@ 3.4 Results of VQA Task
InstructBLIP-7B 0.56 0.65 0.49 0.48 0.36 0.38 0.44 0.43 0.44
Qwen-VL-Chat 0.65 0.82 0.60 0.54 0.51 0.45 0.47 0.51 0.54
LLaVA-V1.5-7B 0.58 0.81 0.54 0.55 0.46 0.45 0.54 0.53 0.53
LLaVA-V1.5-13B 0.70 0.82 0.65 0.60 0.50 0.47 0.58 0.57 0.57
mPLUG-Owl-2 0.51 0.82 0.59 0.55 0.46 0.48 0.47 0.52 0.53
ShareGPT4V-7B 0.58 0.80 0.64 0.54 0.49 0.40 0.51 0.54 0.53
ShareGPT4V-13B 0.67 0.80 0.65 0.56 0.49 0.50 0.6 0.55 0.56
GPT-4V 0.66 0.84 0.73 0.62 0.63 0.63 0.70 0.69 0.67
GPT-4o 0.78 0.91 0.77 0.73 0.75 0.73 0.83 0.77 0.77
Human 0.99 0.96 0.99 0.94 0.96 0.96 0.96 0.93 0.95
Table 5 shows the performance of LVLMs on VQA task. Consistent with results in previous sections, GPT-4o achieves the best performance and there is a performance gap between GPT-4o and GPT-4V. There is also a performance gap between GPT-4 and open-source models. For open-source models, ShareGPT4V-13B and LLaVA-v1.5-13B achieves better performance than other LVLMs based on 7B LLMs, which shows the importance of LLM size to LVLMs on this task. As for LVLMs based on 7B LLMs, the performance of InstructBLIP is the worst and there is also a gap between InstructBLIP and other models. One possible reason is that it only finetunes the Q-former for instruction-tuning as Q-former has limited capacity compared with LLMs. The performance of other models are similar. Consistent with previous findings, reasoning about location is also the easiest for LVLMs and event-related reasoning are more difficult. There is also a large gap between the performance of LVLMs and human beings. Note that the accuracy of Human in Table 5 is calculated based on the responses of five healthy people. They all have obtained a bachelor’s degree and are between the ages of 20 and 30.
@@ 4 Related Work
To better understand emerging capabilities of LVLMs, many scholars have proposed different evaluation benchmarks. LVLM-eHub [24] evaluates LVLMs’ multimodal capabilities across six categories using various publicly available computer vision datasets. MME [10], MMBench [11] and SEED Bench [12] use True/False Questions or Multiple-Choice Questions to evaluate different abilities of LVLMs. MM-VET [25] evaluates LVLMs in terms of their integrated VL capabilities via open-ended questions. Different from them, CogBench mainly focus on high-level cognition evaluation of LVLMs. Though some of them also consider cognition as one of the evaluation dimensions, most images they use can only be used to evaluate limited aspects of reasoning.
@@
Image Captioning is a classical VL task. As model capabilities advance, researchers strive to enhance their ability to describe images in detail. Krause et al. [26] propose Image Paragraph Captioning, tasking models with generating a descriptive paragraph for an image. Recently, some researchers [17, 18, 19, 23] are trying to leveraging the ability of LLMs to generate more detailed image descriptions. None of these tasks consider to evaluate the high-level cognitive ability of models through description. HL dataset [27] requires models to generate high-level captions, but it only considers three aspects (scene, action, rationale). The content of most images in existing datasets does not reach the level of a story. This reveals the need for higher-level datasets like CogBench.
@@
Visual Reasoning is closely related to the cognitive ability of models. VCR [28] tasks models with answering visual questions using commonsense reasoning and justifying their answers. VisualCOMET [29] is a framework of visual commonsense reasoning tasks to predict past, future events, and present intents. Hessel et al. [30] utilize images from VCR and Visual Genome [31] to evaluate the ability of models to perform abductive reasoning. Fu et al. [32] propose a task to identify the time and location of a given image. CURE [33] is proposed to measure both the zero-shot reasoning performance and consistency of VLMs. Similarly, compared CogBench, these tasks considers less kinds of reasoning and CogBench can be seen as the next step of these efforts.
@@ Appendix C Prompt of GPT-based Description Cognition Evaluation
Given a <DESCRIPTION> and some <KEY POINT>s, please tell me if the <DESCRIPTION> explicitly presents the exact or similar semantics of each <KEY POINT>. The following points are required:
1) Instead of reasoning about whether each <KEY POINT> is possibly correct based on the <DESCRIPTION>, you only need to determine whether the <DESCRIPTION> mentions the semantics in the <KEY POINT>. 2) Do not overlook the semantics in the <DESCRIPTION> that are semantically equivalent to the <KEY POINT> but expressed in different ways. For instance, if the <DESCRIPTION> mentions "The woman is playing with her son…", we can tell it successfully includes semantics in the <KEY POINT> "The woman is the mother of the boy." 3) If several possible scenarios are listed using ’or’ at a <KEY POINT>, you only need to determine whether one of these scenarios is mentioned in the <DESCRIPTION>. Assign a score of 0 or 1 to each <KEY POINT>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by models.} <KEY POINT>: 1. {Annotated key point 1.} 2. {Annotated key point 2.} N. {Annotated key point N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <KEY POINT>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 8: Evaluation prompt for GPT-4 of Reasoning types other than [Event Relationship Reasoning].
@@
Figure 9: Evaluation prompt for GPT-4 of [Event Relationship Reasoning].
@@ Appendix D Introduction to Selected LVLMs
Table 6: LVLMs evaluated in this paper.
@@
GPT-4o
  • InstructBLIP [20] builds upon BLIP-2 [34]. It consists of an image encoder, a LLM, and a Query Transformer (Q-Former). During instruction tuning, only the Q-Former is updated. We use “blip2-vicuna-instruct” for testing.
  • Qwen-VL-Chat [21] is a instruction-tuned VL chatbot based on Qwen-VL. Its training process consists of two stages of pre-training and a final stage of instruction fine-tuning. As for architecture, it consists of a vision encoder, a LLM, and position-aware vision-language adapter. We test “Qwen-VL-Chat”.
  • LLaVA v1.5 [22] is an upgraded version of LLaVA [5]. LLaVA connects a vision encoder and LLM for general-purpose visual and language understanding. It is instruction-tuned on the language-image instruction-following data generated by language-only GPT-4 [1]. By using CLIP-ViT-L-336px with an MLP projection and adding academic-task-oriented VQA data with simple response formatting prompts, LLaVA v1.5 achieves better performance. “llava-v1.5-7b” and “llava-v1.5-13b” are tested.
  • mPLUG-Owl-2 [6] mainly comprises a fundamental vision encoder, a visual abstractor, and a language decoder. It also adopts a two-stage training strategy, comprising pre-training and visual instruction tuning. We test “mplug-owl2-llama2-7b”.
  • ShareGPT4V [23] follows the design of LLaVA v1.5. They incorporate a large-scale resource featuring highly descriptive captions into both the pre-training and supervised fine-tuning phases of ShareGPT4V model. We test“ShareGPT4V-7B” and “ShareGPT4V-13B”.
  • GPT-4V [1] is one of the most powerful LVLMs in the world developed by OpenAI. The version of “gpt-4-vision-preview” is tested.
  • GPT-4o [1] is currently OpenAI’s latest and most powerful multimodal model. It is a single model trained end-to-end across text, vision, and audio. The version of “gpt-4o” is tested.
Table 6 shows an overview of the design of different LVLMs.